OpenStudy (anonymous):

Help!! Convert rectangular equation x^2+y^2-3x=0 to polar form. Please Help!! Medal will be given!!1

OpenStudy (ipwnbunnies):

\[x = r*\cos(\theta)\] \[y = r*\sin(\theta)\] \[x^2 + y^2 = r^2\]

OpenStudy (anonymous):

Can you please explain on how to set this problem up what would be r?? @iPwnBunnies

OpenStudy (ipwnbunnies):

'r' represents the distance from the origin, or the point (0,0).

OpenStudy (ipwnbunnies):